OpenStudy (anonymous):

Find a rational and an irrational number between 9.5 and 9.7.

OpenStudy (tkhunny):

I'm going to throw out 9.6 as a possible rational number between 9.5 and 9.7. Too bad - that was the easy one. Now, how about irrational? Hint: 9.5^2 = 90.55 and 9.7^2 = 94.09

OpenStudy (anonymous):

I don't understand this plz give a sampler

OpenStudy (tkhunny):

I did, already. Are there any nice, NOT perfect squares, between 90.55 and 94.09?

OpenStudy (anonymous):

yes 90.00

OpenStudy (tkhunny):

No, that's outside the range I gave.

OpenStudy (anonymous):

/91.00 92.00 93.00

OpenStudy (tkhunny):

That's it. Is \(\sqrt{92}\) in the right range?

OpenStudy (anonymous):

YES

OpenStudy (tkhunny):

Done. Find the decimal equivalent to 100ths. Calculator should do that.